# +SSLIP: Automated Radon-assisted and Rotation-corrected identification of complex HCP slip system activity fields from DIC data

### v1

@@ Abstract
Identification of crystallographic slip in metals and alloys is crucial to understand and improve their mechanical behavior. Recently, a novel slip system identification framework, termed SSLIP (for Slip System-based Local Identification of Plasticity), was introduced to leap from conventional trace-based identification to automated, point-by-point identification that exploits the full deformation kinematics. Using microstructure-correlated deformation data (i.e. sub-micron-scale DIC aligned to EBSD), SSLIP matches the measured in-plane displacement gradient tensor to the kinematics of the optimal combination of multiple slip system activities, at each DIC datapoint. SSLIP was applied and demonstrated to be successful on virtual and experimental case studies of FCC and BCC metals. However, for more advanced and anisotropic HCP crystal structures the complete identification of all slip systems was found to be more challenging, posing limitations on automation and flexibility. Here, we propose a significant extension to the SSLIP framework with the aim of automated slip system identification for complex HCP experiments. The main extensions of the SSLIP method, hereinafter referred to as the +SSLIP method, include (i) a pre-selection of slip systems using a Radon transform, (ii) robustness to measured rigid body rotation by simultaneous identification of the local elastic rotation field, (iii) identification of the two best matching slip systems for each data point, and (iv) a procedure to deal with slip systems with in-plane displacement gradient tensors that cannot be discriminated, yielding the full (HCP) slip system activity maps with all slip systems for each grain. The resulting objective identification method does not rely on, e.g., the Schmid factor to select which slip system is active at each point. We show how slip systems from multiple slip families are successfully identified on virtual and real experiments on a Zn polycrystalline coating. We also provide the full Matlab code for +SSLIP on Github.
@@ 1 Introduction
Understanding the mechanisms of plastic deformation in metals and alloys, governed predominantly by crystallographic slip, is fundamental to advancing material science and engineering. Crystallographic slip, the movement of dislocations across slip planes in specific directions, is intrinsic to the material’s crystal structure and its response to external stresses. The intricacies of this process are encapsulated in various slip systems, dictating the mechanical properties and behavior of materials under load. However, accurately identifying and quantifying slip system activities poses significant fundamental challenges, that are exacerbated by various dislocation mechanisms, such as the complex interplay of multiple slip systems and the presence of diffuse or intersecting slips, as well as the limitations of existing characterization techniques [1, 2, 3, 4, 5, 6].
@@
Over the years, various slip system identification methods have been proposed and applied. Through Scanning Electron Microscopy (SEM), discrete and straight slip bands can be observed and matched to theoretical slip plane traces, based on a measured crystal orientation [7, 8, 9, 10]. Additionally, strain maps acquired through SEM-based Digital Image Correlation (SEM-DIC) can be employed for slip trace analysis [11, 12, 13, 14]. However, in order to obtain a more complete identification, the slip direction also needs to be considered. Therefore, approaches such as the ”Relative Displacement Ratio” (RDR) method [15] and the ”Heaviside DIC method” [16] aim to also extract the slip direction from the SEM-DIC displacement field. The Heaviside DIC method incorporates an additional step function within a DIC subset to deduce the slip trace and direction assuming that the slip is confined to a single slip plane resulting in a clean slip step [16]. On the other hand, the RDR method analyzes regular DIC data by comparing displacement component ratios across a distinct slip trace to theoretical Burgers vector ratios [15]. While these methods have proven effective in various studies [17, 18, 19, 20, 21], their application is mostly limited to scenarios with well-defined slip traces, struggling with diffuse or complex slip, partly pertaining due to the resolution constraints of SEM-DIC.
@@
To address the complexities in slip system identification and the limitations of current methodologies, Vermeij et al. recently introduced an automated identification approach, termed as SSLIP (Slip Systems based Identification of Local Plasticity) [22]. This method leverages the detailed deformation data produced through DIC to facilitate a point-by-point identification of slip system activities across a sample surface. By solving an optimization problem at every SEM-DIC data point, SSLIP matches the measured in-plane kinematics - represented by displacement gradient fields - to a combination of theoretical slip system kinematics derived from Electron Backscatter Diffraction (EBSD) data, after data alignment [4]. This novel approach not only enables the generation of slip system activity fields for each potential slip system but also uncovers the local variations in slip activity, particularly near grain and phase boundaries. Such detailed analysis holds significant promise for enhancing the understanding of plastic deformation mechanisms and allows for more precise comparisons between experiments and theoretical simulations [23, 5, 6, 24], as evidenced by the rapid adoption of the SSLIP method by the research community [25, 26, 27, 28].
@@
However, as discussed by Vermeij et al. [22], the automated and complete application of SSLIP on hexagonal close-packed (HCP) materials is challenging, due to its high anisotropy and wide range of slip families, resulting in a large number of slip systems with (very) similar in-plane kinematics making them linearly dependent in the optimization problem [22]. By analyzing virtual HCP experiments [22], it was found that in such demanding cases, a preselection of slip systems to limit the number of possibly active slip systems was necessary. While it was demonstrated that manual preselection of slip systems works for the virtual HCP experiment (without elastic rotations) [22] and for single and bicrystal Zn experiments [29], automation is warranted in order to sustain statistical investigations of slip activities in multiple grains. The problem is made worse by elastic rotations, which are not included in the SSLIP optimization framework. Even though the apparent deformations due to elastic rotations are tiny, their effect in the presence of a large number of similar slip systems can often be large. As a result, a convincing, robust experimental demonstration of SSLIP on HCP is lacking, especially a demonstration of many HCP grains for which automatic identification is crucial.
@@
Here we propose extend the current SSLIP method to make it a robust automatic slip identification method for HCP (and other complex) crystals. Therefore, we first propose to implement an automate preselection step by means of a Radon transform [30], known for its ability to detect straight features, i.e. slip bands from strain maps [13], and we will demonstrate the efficiency of this automated preselection step in the presence of complex and diffuse slip. Subsequently, from the preselected slip systems, the best-fitting combination of two of the preselected systems will be identified for each individual data point separately, such that all of the preselected systems van be active throughout the complete crystal. This step employs an optimization algorithm that not only seeks the optimal match based on local kinematics but also incorporates a correction for any additional local rotation. This rotation correction is crucial as it accounts for misalignments and grain rotations, that would otherwise impede accurate slip system identification, as explained above. By only allowing two slip systems per point, our approach enhances the robustness of the identification process, while avoiding problems in the iterative solution method due to multiple linear-dependent slip systems. We term this extended methodology as +SSLIP – for Radon-assisted, Rotation-corrected, Slip System based Local Identification of Plasticity – to distinguish it from the original SSLIP method.
@@
The significant advancements made with +SSLIP can be most clearly seen from the challenging virtual HCP experiment initially presented in the original SSLIP paper (Section 7 in [22]). Figure 1(a) displays the strain field of this synthetic experiment, where four slip systems are active from four different slip families. The results from the original full SSLIP identification, shown in Figure 1(b), clearly show that several slip systems were misidentified, as marked by the red crosses. In contrast, Figure 1(c) gives a preview of the results of the new +SSLIP approach, utilizing the Radon transform for automated preselection, showing a nearly perfect identification of all four slip systems, as discussed below.
@@
The paper is structured as follows. In Section 2, the +SSLIP method is outlined in detail and demonstrated on the challenging virtual experiment that is introduced in Figure 1. Next, in Section 3 an experimental SEM-DIC case study on HCP is used to demonstrate the power of the +SSLIP method, for selected HCP grains that show various complicated slip mechanisms, showing a robust identification of slip activity fields while considering no less than 30 slip systems from 5 different slip families. Finally, conclusions are summarized in Section 4.
@@ 2 Methodology
In this section, we outline the computational framework developed to extend the SSLIP methodology to +SSLIP , mainly for the identification of slip system activities within HCP materials. First, the virtual experiment and the original SSLIP method [22] are recapped, after which the new +SSLIP method is introduced.
2.1 Virtual Experiment and Recapitulation of the original SSLIP Method
The virtual experiment, originally designed as the ultimate challenge for the SSLIP methodology [22], incorporates four overlapping slip systems from different slip families that represent the possible complexity of slip activity in an HCP material. The displacement field $\vec{u}$ is generated using the following slip systems: Basal-3 (#3), PrismI-3 (#6), PyrICA-9 (#21), and PyrII-5 (#29). Figure 1(a) already showed the effective strain field $E_{\text{eff}}$, based on the displacement gradient tensor $\mathbf{H}^{exp}$, which is calculated by taking the gradient of the displacement field:
@@ 2.1 Virtual Experiment and Recapitulation of the original SSLIP Method
The slip system’s contribution to the displacement gradient tensor, for a single slip system $\alpha$, is expressed as the product of the slip amplitude $\gamma^{\alpha}$ and the Schmid tensor $\mathbf{P}^{\alpha}$:
$\mathbf{H}^{\alpha}=\gamma^{\alpha}\mathbf{P}^{\alpha}=\gamma^{\alpha}\vec{s}^{\alpha}\otimes\vec{n}^{\alpha},$ (3)
with $\vec{s}^{\alpha}$ and $\vec{n}^{\alpha}$ representing the normalized slip direction and slip plane normal of slip system $\alpha$, respectively. When considering multiple active slip systems, the local theoretical displacement gradient tensor $\mathbf{H}^{theo}$ is the sum of contributions from each system:
@@
The minimization of the sum of absolute values of slip activities is included in order to handle more than 4 slip systems, since there are only 4 measured (known) components of the in-plane 2D displacement gradient tensor. In the original SSLIP paper [22], we showed that this approach works well for a less complicated HCP virtual experiment and also for FCC and BCC real experiments. However, as shown in Figure 1, application to the challenging HCP virtual experiment with 4 overlapping slip systems from different slip families results in erroneous results.
@@
To increase the robustness, the +SSLIP methodology first incorporates a new step for automated preselection of slip systems through the Radon transform, as initially proposed by Hu et al. [13]. An important difference with the method of Hu et al. is that we use the Radon transform only for de-selection of slip systems that clearly cannot be active, while the actual identification is done in the second part of the +SSLIP method, as discussed below. The Radon transform projects a 2D function, in this case the measured effective strain field $E_{eff}$, from Cartesian coordinates ($x$, $y$) to Polar coordinates ($\theta_{pr}$, $x^{\prime}$) [30]:
@@
In this, $\theta_{pr}$ denotes the projection angle and $x^{\prime}$ signifies the radial distance for the given projection angle. This projection into Radon space yields a sinogram that highlights dominant slip orientations as intensity peaks, as illustrated in Figure 2(a-c). In Figure 2(d), the identified slip traces are shown with their corresponding angles, which agree almost perfectly with the measured ones in Figure 2(c).
Figure 2: Application of the Radon transform to the Virtual HCP Experiment. (a) Effective strain field, $E_{\text{eff}}$, with superimposed lines indicating the projection angles, $\theta_{\text{pr}}$, used in the Radon transform. The intersection point, $x_{i}$, marks the center of the local area considered for analysis. (b) The effective strain field with a circular overlay representing the Radon transform’s integration path. The red dot at the center indicates the rotation axis for the Radon transform. (c) The Radon transform output, $R\{E_{\text{eff}}(x,y)\}$, with peaks corresponding to prominent slip band angles in the strain field. (d) Identified slip band orientations superimposed onto the strain field, with color-coded lines representing different slip system, detected by the Radon transform, with corresponding (theoretical) angles.
While this works well on the virtual experiment, we try this now on an experimental strain field to assess its performance. Figure 3(a) shows the strain field of a grain from the dataset of the case study in Section 3. Its Radon transform, shown in Figure 3(b), shows a clear maximum as illustrated with the red ellips, which, as will be shown, does not correspond to the slip band orientations. Therefore, we address the influence of the grain shape by introducing a correction step, similarly to the approach by Hu et al. [13]. A binary mask of the grain shape (Figure 3(c)),
@@
is employed to normalize the Radon transform outputs. The Radon transform of this grain shape function, $g(x,y)$,
[…]
is shown in Figure 3(e). The same Radon maximum values are seen as in Figure 3(d), which confirms that the slip bands are missed. Next, normalization of the peak heights in the original Radon space is performed:
@@
This ensures that the length of integration reflects the actual slip trace length within the grain, as shown in Figure 3(f), wherein the Radon peaks actually correspond to the slip bands. Figure 3(c) shows the strain map again with the identified (Basal) slip trace.
@@
Figure 3: Experimental Grain Analysis Using Normalized Radon transform. (a) Effective strain field within a single grain, $E_{\text{eff}}$, with a magnified inset detailing the grain boundary. (b) Binary mask derived from the strain field to define the grain boundary for normalization. (c) Superimposed identified Basal slip trace on the strain field, conforming to the identified Radon transform peaks. (d) Raw Radon transform, $R\{E_{\text{eff}}(x,y)\}$, displaying the intensity of features as a function of the projection angle, $\theta_{\text{pr}}$, and radial axis, highlighting the maximum values with a red ellipse. (e) Grain shape Radon transform, $R\{g(x,y)\}$, which accounts for the curvature of the grain boundary, showing a similar profile to that of (d). (f) Corrected Radon output with peak indicators and trace lines corresponding to the detected slip bands within the search region, illustrating the precision of slip band orientation identification post grain shape normalization.
This normalization process, using the grain shape function $g(x,y)$, ensures that the length of integration accurately reflects the actual length of slip traces within the grain. The normalized Radon transform, ¯$\overline{\mathcal{R}}\left\{E_{eff}(x,y)\right\}$, isolates the average strain along each projection line, enhancing the precision of slip system orientation identification. Peaks within the normalized sinogram, identified using the peaks2 function in Matlab, represent slip band orientations, thereby automating the trace analysis process and warranting the preselection efficiency and objectivity within the +SSLIP framework.
@@
As a preliminary demonstration, we show in Figure 4 the Radon transform preselected slip systems for all experimental cases considered in this paper. The simple example grain shown in Figure 3 resulted in preselection of only a single slip family, i.e. Basal, in which case it could be argued that further identification with +SSLIP is not required, since a SF based method of, e.g., Hu et al. [13], would perform well here in the case that the slip system with the highest SF within a slip family always activated. However, especially for smaller grains with fewer dislocation sources, this assumption may be questionable. Moreover, Figure 4 shows that all other grains considered in this work show at least 3 preselected slip families (each grain showing at least 6 slip systems). Therefore, it is unreasonable in these cases to rely just on the SF for further selection, especially when the critical resolved shear stresses can be unknown, as is often the case. Additionally, while Figure 4(a,b) show clear slip bands that are visually conforming well to the preselected slip traces, the initial Radon-based preselections of the grains in Figure 4(c-e) are less conclusive. In summary, the Radon transform is a valuable first step, but the SSLIP analysis is required to achieve positive identification of the precise slip system(s).
@@
Figure 4: Demonstration of Radon transform preselection on all grains considered in this work. For each grain, (a-e), the effective strain field $E_{\text{eff}}$ is plotted. The slip traces of the preselected slip systems are drawn as overlay, as described in the legends.
@@
With the Radon transform providing a trace-based preselected set of slip systems, we can now apply SSLIP in a targeted manner. For improved robustness and to avoid erroneous solutions from in-plane linearly dependent slip systems, we follow a refined approach that involves systematically examining every possible combination of two slip systems from the preselected pool, at each datapoint. Essentially, this means conducting SSLIP analysis $N$ times for every datapoint, where $N$ represents the total number of unique two-system combinations from the preselected set:
$N=\frac{N_{radon}!}{2!*(N_{radon}-2)!}.$ (9)
For the current virtual experiment, preselection excludes 16 of the 24 slip systems, only leaving 3 Basal systems, 1 Prismatic system, 2 PyrICA systems and 2 PyrII systems to be included ($N_{radon}$ = 8). Thus, according to Eq. 9, this results in $N=28$ combinations to be analyzed by SSLIP. Below, we will detail the criteria and process for selecting the optimal combination for each datapoint, ensuring a precise and comprehensive identification of active slip systems.

### v2

@@ Abstract
Identification of crystallographic slip in metals and alloys is crucial to understand and improve their mechanical behavior. Recently, a novel slip system identification framework, termed SSLIP (for Slip Systems based Local Identification of Plasticity), was introduced to leap from conventional trace-based identification to automated, point-by-point identification, exploiting the full deformation kinematics. Using sub-micron-scale Digital Image Correlation (DIC) deformation fields aligned to Electron Backscatter Diffraction (EBSD) data, SSLIP matches the measured in-plane displacement gradient tensor to the kinematics of the optimal combination of multiple slip system activities, at each DIC datapoint. SSLIP was demonstrated to be successful on virtual and experimental case studies of FCC and BCC metals. However, for more challenging HCP crystal structures, the complete identification of all slip systems was found to be more challenging, posing limitations on automation and flexibility. To extend the capabilities of SSLIP, we propose an extended framework, hereinafter referred to as the +SSLIP method, which includes (i) a preselection of slip systems using a Radon transform, (ii) robustness to measured rigid body rotation by simultaneous identification of the local rotation field, (iii) identification of the two best matching slip systems for each data point, and (iv) a procedure to determine groups of slip systems with in-plane displacement gradient tensors that cannot be discriminated. This procedure yields the full (HCP) slip system activity maps for every slip system in each grain. The resulting objective identification method does not rely on the Schmid factor to select which slip system is active at each point. We show how slip systems from multiple slip families are successfully identified on virtual and real experiments on a Zn polycrystalline coating. https://doi.org/10.1111/str.70000
@@ 1 Introduction
Understanding the mechanisms of plastic deformation in metals and alloys, governed predominantly by crystallographic slip, is fundamental to advancing materials science and engineering. Crystallographic slip, comprising the movement of dislocations across slip planes in specific crystallographic slip directions, is a crucial plastic deformation mechanism that governs how metals respond to external stresses. The combination of a certain slip plane and slip direction forms a slip system and is used to describe the kinematics and geometry of plastic slip. However, accurately identifying and quantifying slip system activities in experiments poses significant fundamental challenges, that are exacerbated by various complex phenomena, such as the complex interplay of multiple slip systems and the presence of diffuse or intersecting slip bands, which are hard to measure with existing characterization techniques [1, 2, 3, 4, 5, 6].
@@
Over the years, various slip system identification methods have been proposed and applied. Through Scanning Electron Microscopy (SEM), discrete and straight slip bands can be observed and matched to theoretical slip plane traces, based on measured crystal orientations through Electron Backscatter Diffraction (EBSD) [7, 8, 9, 10]. Additionally, strain maps acquired through SEM-based Digital Image Correlation (SEM-DIC) can be employed for slip trace analysis [11, 12, 13, 14]. However, in order to obtain a more complete identification, the slip direction also needs to be considered. Therefore, approaches such as the ”Relative Displacement Ratio” (RDR) method [15] and the ”Heaviside DIC method” [16] aim to also extract the slip direction from the SEM-DIC displacement field. The Heaviside DIC method incorporates an additional step function within a DIC subset to deduce the slip trace and direction, assuming that slip is confined to a single slip plane resulting in a clean slip step [16]. On the other hand, the RDR method analyzes regular DIC data by comparing displacement component ratios across a distinct slip trace to theoretical Burgers vector component ratios [15]. While these methods have proven effective in various studies [17, 18, 19, 20, 21], their application is mostly limited to scenarios with well-defined slip traces, and they struggle with diffuse or complex slip, partly pertaining due to the resolution constraints of SEM-DIC.
@@
To address the complexities of slip system identification and the limitations of current methodologies, Vermeij et al. recently introduced an automated identification approach, termed as SSLIP (for Slip Systems based Local Identification of Plasticity) [22]. This method leverages the detailed deformation data produced through DIC to facilitate a point-by-point identification of slip system activities across a sample surface. By solving an optimization problem at every SEM-DIC data point, SSLIP matches the measured in-plane kinematics - represented by displacement gradient tensor fields - to a combination of theoretical slip system kinematics derived from EBSD data after data alignment [4]. This novel approach not only enables the identification of slip system activity fields for each potential slip system, but also discloses the local variations in slip activity, particularly near grain and phase boundaries. Such detailed analysis holds significant promise for enhancing the understanding of plastic deformation mechanisms and allows for more reliable comparisons between experiments and theoretical simulations [23, 5, 6, 24], as evidenced by the rapid adoption of the SSLIP method by the research community [25, 26, 27, 28].
@@
However, as discussed by Vermeij et al. [22], the automated and complete application of SSLIP on Hexagonal Close-Packed (HCP) materials is challenging due to its high anisotropy that enables a wide range of slip families. This results in a large number of slip systems with (very) similar in-plane kinematics, making them linearly dependent in the optimization problem [22]. By analyzing virtual HCP experiments, it was found that in such demanding cases, a preselection of slip systems to limit the number of possibly active slip systems was necessary to achieve a correct identification. While it was demonstrated that manual preselection of slip systems works for the virtual HCP experiment (without elastic rotations) and for single and bicrystal Zn experiments [29], automation is warranted in order to sustain statistical investigations of slip activities in multiple grains. The issue is exacerbated by elastic rotations, which are not accounted for in the SSLIP optimization framework. Even though the apparent deformations due to elastic rotations are tiny, their effect in the presence of a large number of similar slip systems can often be large. As a result, a convincing, robust experimental demonstration of SSLIP on HCP is lacking, especially a demonstration of many HCP grains, for which automatic identification is crucial.
@@
Here we extend the current SSLIP method to make it a robust and automatic slip identification method for HCP (and other complex) crystals. Therefore, we first propose to implement an automated preselection step by means of a Radon transform [30], known for its ability to detect straight features, i.e. slip bands from strain maps [13]. We will demonstrate the efficiency of this automated preselection step in the presence of complex and diffuse slip. Subsequently, from the preselected slip systems, the best-fitting combination of two of the preselected systems will be identified for each individual data point separately. This step employs an optimization algorithm that not only seeks the optimal match based on the local kinematics but also incorporates a correction for any additional local rotation. This rotation correction is crucial as it accounts for rotational misalignment of the sample and grain rotations, that may otherwise impede accurate slip system identification, as explained above. By only allowing two slip systems per point, our approach enhances the robustness of the identification process, while avoiding problems in the iterative solution method due to multiple linear-dependent slip systems. We expect this to be a reasonable restriction, as it is unlikely that more than two slip systems are active in a single data point. Finally, we also identify groups of nondiscriminatory slip systems, of which the in-plane kinematics cannot be discerned. We term this extended methodology as +SSLIP to distinguish it from the original SSLIP method.
@@
As a preview, we show the significant advancements made with +SSLIP on the challenging virtual HCP experiment presented in the original SSLIP paper (Section 7 in [22]). Figure 1(a) displays the strain field of this synthetic experiment, where four slip systems are active from four different slip families. The results from the original full SSLIP identification (without preselection), shown in Figure 1(b), clearly show that several slip systems were misidentified, as marked by the red crosses. In contrast, Figure 1(c) gives a preview of the results of the new +SSLIP approach, utilizing the Radon transform for automated preselection, showing a nearly perfect identification of all four slip systems, as discussed below.
@@
The paper is structured as follows. In Section 2, the +SSLIP method is outlined in detail and demonstrated on the challenging virtual experiment that is introduced in Figure 1. Next, in Section 3, an experimental SEM-DIC case study of HCP is used to demonstrate the strength of the +SSLIP method for selected HCP grains that show various complicated slip mechanisms, showing a robust identification of slip activity fields while considering no less than 30 slip systems from 5 different slip families. Finally, conclusions are summarized in Section 4.
@@ 2 Methodology
In this section, we outline the framework developed to extend the SSLIP methodology to +SSLIP , focusing on the identification of slip system activities within HCP materials. First, the virtual experiment and the original SSLIP method [22] are reiterated, followed by the introduction of the new +SSLIP method.
2.1 Virtual Experiment and Reiteration of the Original SSLIP Method
The virtual experiment, originally designed as the ultimate challenge for the SSLIP methodology [22], incorporates four overlapping slip systems from different slip families that represent the possible complexity of slip activity in an HCP material. The displacement field $\vec{u}$ is generated using the following slip systems: Basal-3 (#3), PrismI-3 (#6), PyrICA-9 (#21), and PyrII-5 (#29). As an example, ”PrismI-3 (#6)” refers to the third system of the first order Prismatic slip system family, which is the sixth slip system overall. Figure 1(a) already showed the effective strain field $E_{\text{eff}}$, based on the displacement gradient tensor $\mathbf{H}^{exp}$, which is calculated by taking the gradient of the displacement field:
@@ 2.1 Virtual Experiment and Reiteration of the Original SSLIP Method
The slip system’s contribution to the displacement gradient tensor, for a single slip system $\alpha$, is expressed as the product of the slip magnitude $\gamma^{\alpha}$ and the Schmid tensor $\mathbf{P}^{\alpha}$:
$\mathbf{H}^{\alpha}=\gamma^{\alpha}\mathbf{P}^{\alpha}=\gamma^{\alpha}\vec{s}^{\alpha}\otimes\vec{n}^{\alpha}.$ (3)
Here, $\vec{s}^{\alpha}$ and $\vec{n}^{\alpha}$ represent the normalized slip direction and slip plane normal of slip system $\alpha$, respectively. When considering multiple active slip systems, the local theoretical displacement gradient tensor $\mathbf{H}^{theo}$ is the sum of the contributions of each system:
@@
The minimization of the sum of absolute values of slip activities is included in order to handle more than four slip systems, since there are only four measured (known) components of the in-plane 2D displacement gradient tensor. In the original SSLIP paper [22], we showed that this approach works well for a less complicated HCP virtual experiment and also for FCC and BCC real experiments. However, as shown in Figure 1, application to the challenging HCP virtual experiment with 4 overlapping slip systems from different slip families results in erroneous results.
@@
To increase the robustness, the +SSLIP methodology first incorporates a new step for automated preselection of slip systems through the Radon transform, as proposed by Hu et al. [13]. An important difference with the method of Hu et al. is that we use the Radon transform only for deselection of slip systems that clearly cannot be active, while the actual identification is performed in the second part of the +SSLIP method, as discussed below. The Radon transform integrates a 2D function, in this case the measured effective strain field $E_{eff}$, along a single line (limited to $x^{\prime}=x\cos\theta_{pr}-y\sin\theta_{pr}$ by usage of the Dirac delta function $\delta$) and projects that value into Radon space ($\mathcal{R}\{E_{eff}(x,y)\}$), which is spanned by the variable $\theta_{pr}$ and $x^{\prime}$ [30, 13]:
@@
Here, $\theta_{pr}$ is the positive angle of the line with respect to the original y-axis, while $x^{\prime}$ denotes the smallest distance between the line and the origin of the cartesian coordinate system. Line integrals are performed along all possible lines across the given data set, which yields a sinogram that highlights dominant slip bands as high intensity peaks, as shown in Figure 2(a,b). The measured slip band angles are then matched to theoretical slip trace orientations, using a threshold of $\pm$10 . In Figure 2(c), the identified slip traces are shown with their corresponding angles, which agree perfectly with the measured ones in Figure 2(c).
Figure 2: Application of the Radon transform to the Virtual HCP Experiment. (a) Effective strain field, $E_{\text{eff}}$, with superimposed lines indicating the projection angles, $\theta_{\text{pr}}$, used in the Radon transform. The intersection point $x_{i}$ marks the center of the local area considered for analysis. The red dot at the center indicates the rotation axis for the Radon transform. (b) The Radon transform output, $R\{E_{\text{eff}}(x,y)\}$, with peaks corresponding to prominent slip band angles in the strain field. (c) Identified slip trace orientations superimposed onto the strain field, with color-coded lines representing different slip system, detected by the Radon transform, with corresponding theoretical angles.
While this works well on the virtual experiment, we try this now on an experimental strain field to assess its performance. In order to avoid the influence of grain boundary deformation, we first remove a band of 5 pixels along the grain boundary. Figure 3(a) shows the strain field of a grain from the dataset of the case study in Section 3. Its Radon transform, shown in Figure 3(d), shows a clear maximum as illustrated with the red ellips, which, as will be shown, does not correspond to the actual slip band orientations. Therefore, we address the influence of the grain shape by introducing a correction step, similar to the approach by Hu et al. [13]. A binary mask of the grain shape (Figure 3(c)),
@@
is employed to normalize the Radon transform outputs. The Radon transform of this grain shape function $g(x,y)$ ($\mathcal{R}\{g(x,y)\}$) is shown in Figure 3(e). The same Radon maximum values are seen as in Figure 3(d), which confirms that the slip bands are not identified. Next, normalization of the peak heights in the original Radon space is performed:
@@
This ensures that the length of integration reflects the actual slip trace length within the grain, as shown in Figure 3(f), wherein the Radon peaks actually correspond to the slip bands.
Slip band orientations are identified by finding peaks in the normalized Radon sinogram (Figure 3(f)). We employ the peaks2 function iteratively in Matlab [32]. peaks2 operates by comparing each data point to its eight immediate neighbors in a vectorized manner, ensuring efficient execution. A point is considered a peak if it is strictly greater than all its surrounding neighbors. Relevant parameters are summarized in Table 1. Starting with a minimum peak height of 1 (maximum theoretically possible), the minimum peak height is lowered in steps of $IncPeakHeight$ until the desired number of peaks ($NmaxPeaks$) are identified, or until the overall minimum peak height ($MinPeakHeight$) is reached. The peaks need to be spaced at least $MinPeakDistance$ pixels. If multiple peaks are identified within that range, then the one with the highest intensity is chosen as the only peak. The slip band angles corresponding to the identified peaks are then compared to theoretical slip trace angles (from EBSD) and must fall within a threshold of $OrientThresh$ to be considered a potential slip system. The result is shown in Figure 3(f), in which two peaks are identified. Note that the next highest peaks have the same angle, meaning that the net result would be the same, which we found is often the case. Moreover, missing several minor peaks of one slip system is not an issue, since the purpose is only to identify the presence of active slip systems using the Radon, not to identify their location or amount of active bands. For both reasons, the precise peak finding parameters are not very important, as far as we can tell. Finally, Figure 3(c) shows the strain map again with the identified (Basal) slip trace.
Table 1: Parameters for peak identification in Radon transform
Parameter Value Unit
$MinPeakHeight$ 0.8 [-]
$NmaxPeaks$ 5 [-]
$MinPeakDistance$ 50 [pix]
$IncPeakHeight$ 0.025 [-]
$OrientThresh$ 10 []
@@
Figure 3: Experimental grain analysis Using Normalized Radon transform. (a) Effective strain field within a single grain, $E_{\text{eff}}$, with a magnified inset detailing the grain boundary, around which data is removed. (b) Binary mask derived from the strain field to define the grain boundary for normalization. (c) Superimposed identified Basal slip trace on the strain field, conforming to the identified Radon transform peaks. (d) Raw Radon transform, $R\{E_{\text{eff}}(x,y)\}$, highlighting the maximum values with a red ellipse. (e) Grain shape Radon transform, $R\{g(x,y)\}$, which accounts for the curvature of the grain boundary, showing a similar profile to that of (d). (f) Corrected Radon output with peak indicators (Magnified in the inset) and trace lines corresponding to the detected slip bands within the search region, illustrating the precision of slip band orientation identification post grain shape normalization.
@@
As a preliminary demonstration, we show in Figure 4 the Radon transform preselected slip systems for all experimental cases considered in this paper. The simple example grain shown in Figure 3 resulted in preselection of only one single slip family, i.e. Basal slip, see Figure 4(f). In such cases, it could be argued that further identification with +SSLIP is not required, since a Schmid Factor (SF) based method of, e.g., Hu et al. [13], would perform well here in the case that the slip system with the highest SF within a slip family is always activated. However, the SF is known to be unreliable for polycrystalline microstructures [15]. Additionally, for smaller grains with fewer dislocation sources, the activation of slip systems is more stochastic, making the SF less useful. Moreover, Figure 4(a-e) shows that all other grains considered in this work show at least 3 preselected slip families (at least 6 slip systems for each grain). Therefore, it is unreasonable in these cases to rely just on the SF for further selection, especially when the critical resolved shear stresses can be unknown and different for the different slip families, as is often the case. Additionally, while Figure 4(a,b) show clear slip bands that are visually conforming well to the preselected slip traces, the initial Radon-based preselections of the grains in Figure 4(c-e) are less conclusive. In summary, the Radon transform is a valuable first step, but the SSLIP analysis is required to achieve positive identification of the precise slip systems and their activity fields.
@@
Figure 4: Demonstration of Radon transform preselection on all grains considered in this work. For each grain, (a-f), the effective strain field $E_{\text{eff}}$ is plotted. The slip traces of the preselected slip systems are drawn as overlay, as described in the legends. Note that the PyrIA slip traces are not visible since they overlap with the PyrICA slip traces.
@@
With the Radon transform providing a trace-based preselected set of slip systems, we can now apply SSLIP in a targeted manner. For improved robustness and to avoid erroneous solutions from in-plane linearly dependent slip systems, we follow a refined approach that involves systematically examining every possible combination of two slip systems from the Radon transform preselected pool (of $N_{radon}$ slip systems), at each datapoint. Essentially, this means conducting the SSLIP analysis $N_{duo}$ times for every datapoint, where $N_{duo}$ represents the total number of unique two-system combinations (i.e. duos) from the preselected set:
$N_{duo}=\frac{N_{radon}!}{2!*(N_{radon}-2)!}.$ (8)
For the virtual experiment introduced earlier, preselection excludes 16 slip systems, only leaving 3 Basal systems, 1 Prismatic system, 2 PyrICA systems and 2 PyrII systems to be included ($N_{radon}$ = 8). Thus, according to Eq. 8, this results in $N_{duo}=28$ combinations to be analyzed by SSLIP. Below, we will detail the criteria and processes for selecting the optimal combination for each datapoint, ensuring a precise and comprehensive identification of active slip systems.